Answer the following and give complete solutions.

1. Find the constant a so that phi(x)=x(e^2x) is a solution to y'+ay-e^(2x)=0
2. Find the constant a so that e^xy + y = x-1 is an implicit solution to dy/dx = e^-xy +ay / e^-xy + x.
1
Expert's answer
2020-05-31T19:01:54-0400

A)Givenϕ (x) = xe2x


"\phi '= e^{2x}(1 + 2x)"

If  ϕ(x) is a solution to the given equation then it satisfies the equation,

"e^{2x}(1+2x)+axe^{2x}-e^{2x}=0 \\
1+2x+ax-1=0 \\
2x+ax=0 \\
a=-2"


B)

Given y(ex+1)=x-1

Differentiating with respect to x

dy ( ex+1) + y(ex) dx = dx

dy(ex+1) = dx ( 1-yex)

"\frac{dy}{dx} = \frac{1-ye^{x}}{e^{x} +1}"


Putting this value in the given differential equation

"\frac{1-ye^{x}}{e^{x} +1}=e^{-x}y+\frac{ay}{e^{-x}y}+x\\"


"\frac{1-ye^{x}}{e^{x} +1}-e^{-x}y-x" ="\frac{a}{e^{-x}}"

 

"\frac{1-ye^{-x}-y(1+e^x)-x(1+e^x)}{1+e^x}= ae^x \\

\\

\frac{1-ye^{-x}-(1+e^x)(x+y)}{e^x(1+e^x)}= a"